Home Decor Pty Ltd is considering investing in a new machine to assemble its new product of high quality speakers. The machine is estimated to cost $124000 which can last for 5 years before it becomes unreliable and can be sold for scrap at $12,000. The project is estimated to bring in additional $40,000 net cash inflow annually with an annual 2% growth . The net cash flow in year 5 also includes the scrap value. The company plans to fund the purchase of the new machine using a bank loan with an interest rate of 10%. A. What is the ARR of the project?(Case sensitive. Type in 20.00 (two decimal places) for 20%.) B. How long is the payback period for this project? (Case sensitive. Type in 7.00 (two decimal places) for 7 years.) C. What is the NPV for this project? $ D. What is the IRR for this project? %
Answer #1
Annual depreciation (124000-12000)/5 22400
Year 1 2 3 4 5
annual cash flow =cash flow in year 1*(1+g)^n growth = 2% 40000 40800 41616 42448.32 43297.2864
average accounting income = annual cash flow-annual depreciation 17600 18400 19216 20048.32 20897.2864
average cash flow = sum of annual cash flow/n 19232.32128
Accounting rate of return = average accounting income/initial investment 19232.32/124000 15.51%
2-
Year 1 2 3 4 5
annual cash flow =cash flow in year 1*(1+g)^n growth = 2% 40000 40800 41616 42448.32 43297.2864
scrap value of equipment 12000
net operating annual cash flow -124000 40000 40800 41616 42448.32 55297.2864
cumulative cash flow 40000 80800 122416 1584
amount to be recovered in year 4
Payback period in years = year before final year of recovery+(amount to be recovered in year 4/net operating annual cash flow) 3+(1584/42448.32) 3.04
3-
Year 0 1 2 3 4 5
net operating annual cash flow -124000 40000 40800 41616 42448.32 55297.2864
present value factor at 10% = 1/(1+r)^n r =10% 1 0.909090909 0.826446281 0.7513148 0.68301346 0.62092132
present value of net operating cash flow= net operating cash flow*present value factor at 10% -124000 36363.63636 33719.00826 31266.7168 28992.7737 34335.2642
net present value = sum of present value of net operating cash flow 40677.40
4-
Year 0 1 2 3 4 5
net operating annual cash flow -124000 40000 40800 41616 42448.32 55297.2864
IRR = Using IRR function in MS excel IRR(G3734:L3734) 21.54%

. Co New Microsoft Office Excel Worksheet - Microsoft Excel Insert Page Layout Formulas Data Review View Home Cut Dr Percenta